# What’s under the hood: Investigating Automatic Metrics on Meeting Summarization

### v1

@@ 4.3 Metrics
For aligning human annotation with automatic metrics, we employ a suite of metrics spanning count-based, model-based, and QA-based methodologies, chosen based on their prevalent use in meeting and dialogue evaluation Gao and Wan (2022).
4.3.1 Count-based
@@
4.3.2 Model-based
BERTScore Zhang et al. (2020b) utilizes a pre-trained BERT model to measure the contextual similarity between the generated and reference texts, reflecting both semantic and syntactic similarity. We report the rescaled  [^3] https://github.com/Tiiiger/bert score/blob/master/journal/rescale baseline.md BERTScore-F score. Perplexity (PPL) measures the uncertainty of a language model in predicting each word in a given text, with lower scores denoting better predictions. It evaluates the model’s grasp of language structure and word occurrence patterns, thereby evaluating the quality and fluency of utterances. We use GPT-2 Radford et al. (2019) as the language model. BLANC Lita et al. (2005) measures how well a generated summary aids a pre-trained language model in performing a language understanding task on the original document. The score reflects the informativeness of the generated text. LENS Maddela et al. (2023) is a trainable metric which aims to reflect how closely the generated text aligns with human references in terms of content and style.
4.3.3 QA-based
@@
QuestEval Scialom et al. (2021) combines scores obtained through the FEQA and SummaQA metrics. It employs a question-answering model to answer questions formed from the reference text (SummaQA) or the generated summary (FEQA), extracting answers from the opposite source. QuestEval evaluates factual accuracy, coherence, informativeness, and relevance.
4.4 Human Annotation
We adapted proven methodologies for a thorough annotation process Zhang et al. (2023). Four graduate students, aged 24 to 28, from diverse academic backgrounds (i.e., computer science, psychology, communication science), all well-versed in English and familiar with meeting summarization, were involved.
@@ 4.4 Human Annotation
From the QMSum general-summary test set, we chose the 35 general-summary samples, each containing meeting transcripts, gold standard summaries, and model-generated summaries, resulting in a total of 175 distinct samples for annotation. Annotators identified challenges Section 3.1 in the transcript and errors Section 3.2 in the generated summaries. The evaluation used yes/no questions, such as: "Does the given summary omit crucial information or provide insufficient detail about salient points?" corresponding to the missing information error. They further rated the frequency of challenges and the impact of errors for the LED model, using Likert scales from rare occurrence or minimal impact (1) to frequent presence or high impact (5).
To ensure reliable and consistent annotation, we implemented several measures. We assessed inter-annotator agreement using Krippendorff’s alpha, achieving an average of 0.81 (see Tables 2 and 3). Before the annotation task, a pilot test was administered as training for our annotators and to refine our guidelines and point out new error types at the same time. Using the validation dataset, the annotators practiced on QMSum summaries produced by the SLED+BART model Ivgi et al. (2022). For the actual annotation task, we further used gold summaries that the annotators should be able to evaluate correctly, otherwise, their annotation would be rejected. Regular review meetings maintained consistency and alignment during the annotation process, and for ongoing quality control, annotators highlighted summary segments where they detected errors. Discrepancies were discussed, with an expert annotator on standby for complex issues.
@@ 5.1 Linking Challenges and Error Types
Errors Encoder-Decoder Autoregressive
@@
incorrect reasoning speaker dynamics, implicits context
@@
corefernce*, contextual turns
@@
Table 4: The linkage between challenges and errors. No asterisk indicates low correlation, * signifies mid correlation, and ** denotes high correlation.
We analyzed the relationship between challenges and frequently observed errors for encoder-decoder and autoregressive architectures, using Point-biserial correlation based on the human annotations from Section 4.4. The results are presented in Table 4. For encoder-decoder models, errors such as incoherence, structural disorganization, and redundancy arise from most challenges. These errors are strongly linked with challenges like spoken language and speaker dynamics, suggesting encoder-decoder models struggle with maintaining coherence in these complex challenges. Conversely, wrong references, linguistic inaccuracy, and hallucination errors show limited associations with the examined challenges. Specifically, the limited association of hallucination reinforces current understandings, suggesting that the underlying causes remain elusive Maynez et al. (2020). Notably, the implicit context challenge works as a proxy for hallucination. The minimal correlation of the wrong reference error, which rarely occurs due to the sentence structure in generated summaries, and the linguistic inaccuracy error indicate that challenges such as coreference are well-handled by these models.
Regarding specific challenges, discourse structure mildly correlates with the missing information error, suggesting that while models navigate the broader meeting structure well, they occasionally miss details within distinct phases. The low information density challenge weakly correlates with the incorrect reasoning error, revealing occasional difficulties in salient meeting details extraction. However, the contextual turn-taking challenge aligns more with errors like redundancy and incoherence, revealing issues in capturing the dynamics on a granular level within the different meeting phases.
For autoregressive models, patterns differ. Incorrect reasoning is now the rarest observed error. It shows the weakest correlation, while wrong references, redundancy, and linguistic inaccuracy errors are the most prevalent, aligning with challenges like low information and spoken language. This reflects the model’s design, as it presents a shift in redundancy and linguistic inaccuracies compared to encoder-decoder models, with redundancies manifesting as introductory sentences that are then repetitively echoed in subsequent sentences. Linguistic inaccuracies, on the other hand, often emerge as contextually ambiguous terms. Equally prominent but stronger correlated are now the missing information and structural disorganization errors, showing links to the coreference challenge and emphasizing models’ tendencies to list topics without proper context.
Notably, some negative correlations also emerge, suggesting specific challenges might decrease the likelihood of certain errors, such as for low information density challenge and missing information error. These novel observations in the field warrant further exploration.
@@ 5.2 Correlation of Automatic Metrics and Human Annotation
Though numerous and frequently applied, existing automatic metrics aren’t tailored to the intricacies of meeting summarization. We analyze their reactions to the diverse error types observable in meeting summaries by answering the following three research questions.
@@
Table 5: Point-biserial correlation between automatic metrics and annotated errors. * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01. The top row shows abbreviated error types from Section 3.2 in order. A negative correlation indicates worsening metric scores with increasing error occurrence. The three highest absolute values are highlighted in bold. We present F-scores (F) for BERTScore and QuestEval, with BERTScores being further rescaled.
@@
In Table 5, we present Point-biserial correlations between automatic metrics and expert-annotated errors, specifically for LED model-generated summaries due to their pronounced error variance. Our analysis highlights some overarching patterns: No metric consistently correlates highly with all error types, underscoring the complexity of meeting summarization and the absence of a universal evaluative metric that captures human judgment well Gao and Wan (2022). However, some metrics show trends aligning with human judgment through significant negative scores. Though not designed for structural coherence, several metrics show significant negative correlations with structural disorganization errors. This indicates that such temporal and logical disorganization breaks the n-gram sequences and semantic flow. This trend can be observed as the score of ROUGE-1 is vastly more significant than the ROUGE-2 and ROUGE-L scores, aligning with observations in dialogue summarization Gao and Wan (2022). In particular, the gaps, disjoint narratives, and incorrect statements from errors such as missing information, incoherence, and incorrect reasoning influence metric scores, as these events do not align with n-grames in the reference or shift meaning, thereby affecting the semantic alignment Lin (2004). Errors such as redundancy, wrong references, and hallucination remain less frequently detected in metric evaluation scores. This may stem from the fact that summaries with these errors, depending on how many words were altered, can still closely match the reference regarding n-gram overlap and semantic similarity.
@@
As anticipated, count-based metrics directly compare generated and reference texts and are acutely responsive to missing information Lin (2004). ROUGE stands out, with ROUGE-1, in particular, correlating significantly with information omissions, aligning with prior dialogue summarization studies Gao and Wan (2022). Similarly, BLEU identifies summaries with critical point omissions. Despite not being designed for it, count-based metrics significantly correlate with the structural disorganization error, possibly arising from disrupted n-gram sequences due to reorganized content Banerjee and Lavie (2005). Moreover, ROUGE and Meteor are potential indicators for incorrect reasoning and incoherence.
@@
Among model-based metrics, BERTScore exhibits sensitivity to missing information, an anticipated trend possibly stemming from the score reflecting this in the semantic and contextual differences between candidate and gold summary in its score Zhang et al. (2020b). However, its correlation with structural disorganization is less direct, hinting at the influence of sentence alignment and structural coherence. BLANC, designed for evaluating coherence and fluency, naturally correlates negatively with incoherence errors. The text simplification metric, LENS, effectively captures redundancy errors. Model-based metrics generally present milder correlations, possibly due to discrepancies between training data and meeting contexts Gao and Wan (2022). While they seem to not align as well with human judgment as count-based metrics, they do offer insights into specific errors not captured by count-based metrics.
@@
The correlation table shows that no single metric predominantly captures all error types, with most correlations being weak to moderate. Examining Table 5, a combination of metrics like ROUGE, BLANC, and LENS could serve as a proxy for various errors due to their notable moderate correlations. Still, this approach warrants additional evaluation and score weighting.
@@
Table 6: Spearman correlation between metrics and annotated error impacts. * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01. The top row lists abbreviated error types from Section 3.2 in sequence. A negative correlation implies declining metric scores with rising error instances. The three most pronounced absolute values are emphasized in bold. F-scores (F) are given for BERTScore and QuestEval, with BERTScores being further rescaled.
@@
We categorize "masking" in two ways: indifference to an error, seen as a near-zero correlation, or positive reinforcement of a mistake, seen as a positive correlation. Count-based metrics predominantly show near-zero or negative correlations with errors. This indicates that while they might not favour errors, they do not sufficiently penalize specific mistakes. Such behaviour is expected Saadany and Orasan (2021); Akter et al. (2022), especially for errors that these metrics weren’t explicitly designed to detect or for which they can only act as a proxy.
In contrast, while some metrics show their known struggle with specific error types through near-zero scores (e.g., BERTScore and linguistic inaccuracy Hanna and Bojar (2021)), several model-based and QA-based metrics occasionally manifest positive correlations. For instance, QuestEval rewards errors like missing information and hallucination, a counterintuitive outcome given its designed focus on factual accuracy. The finding indicates that hallucinated details subtly bypass metric detection, likely due to the generated questions lacking the rigour to spot hallucinations, which can be coupled to the gap between the pre-training and meeting data Gao and Wan (2022). Similarly, BERTScore, emphasizing semantic and syntactic similarities, may reward redundant yet correct details Hanna and Bojar (2021). In contrast, Perplexity favours incorrect references if they preserve linguistic coherence and fluency, thus aligning with the language model’s expectations and yielding a lower (better) score. Notably, LENS significantly correlates with the structural disorganization error. LENS, although designed for text simplification and primarily emphasizing word-level and intra-sentence simplification Maddela et al. (2023), appears to struggle with capturing broader logical and temporal structures, resulting in its significant correlation with the structural disorganization error.
@@
To further scrutinize the relationship between metric scores and error severity, we examined how these scores fluctuate by the impact of errors. The Spearman correlation trends are depicted in Table 6. Across different models and error severities, most correlations are negligible to weak, and many lack statistical significance, emphasizing the limitations of current metrics in discerning error severity in meeting summarization. BLEU and QuestEval display significant positive correlations with hallucination errors, underscoring their vulnerability to hallucinated content. While some metrics, such as BERTScore and LENS, showed sensitivity to missing information in RQ1, their precision in assessing the error severity remains limited. This might stem from the gap between the pre-training data and meeting transcript for the three metrics utilising language models, as the meeting transcript was provided as part of the input Gao and Wan (2022). LENS, being closely aligned with text simplification, penalizes redundancy, while Perplexity is particularly sensitive to wrong references and linguistic inaccuracies. These observations underscore that, while not explicitly designed for certain errors, some metrics are intrinsically responsive to them.
@@ 6 Related Work
Research on meeting summarization has typically addressed challenges and error types in isolation, leaving their interrelation largely unexplored. For the parent-domain dialogue summarization, challenges have been collected Chen and Yang (2020b). While we adapted these to our domain and corpus, we found them lacking the detail required for meeting contexts. Previous works, especially by Tang et al. (2022); Wang et al. (2022b), laid the groundwork for our error typologies. This foundation was enhanced with insights from diverse studies and feedback from our annotators. Further, while Chen and Yang (2020b) correlate challenges and errors in dialogue summarization, their findings didn’t entirely align with our needs, though their annotation approach informed ours. Although automatic metrics for meeting summaries are underexplored, similar studies exist for text summarization using the CNN/Dailymail dataset Fabbri et al. (2021) and for dialogue summarization via the SAMSum dataset Gao and Wan (2022). We built upon these works by selectively adopting their metrics and augmenting them with measures commonly reported in recent meetings and dialogue summarization studies.
@@ 7 Conclusion
We’ve crafted a resource suite dedicated to meeting summarization evaluation, encompassing domain-specific challenges, linked error types, their predictive correlations for a given model architecture with a known set of challenges, and expert annotations of a QMSum subset. Our analysis highlights the constraints and misalignments of current metrics with human judgment in discerning error nuances. Metrics working well for other summarization tasks either do not react to errors or cannot reflect the impact on the quality of their score. A composite metric might be more effective, although its formulation warrants further research. Recent advancements in the community have highlighted the potential of LLM-based metrics, particularly with chain-of-thought or tree-of-thought prompting techniques. Utilizing an LLM like GPT-4 or Llama 2, prompted with our detailed annotator guidelines and supplemented by a few examples, presents a promising approach. This method could be especially effective in detecting complex errors like structural disorganization and incorrect reasoning. In the future, we plan to augment the entire QMSum dataset with our annotations. Drawing inspiration from the Stanford Context Word Similarity dataset’s impact on word sense ambiguity Huang et al. (2012), we see an enriched QMSum offering deeper insights into potential challenges for new techniques. Further, investigating additional datasets (e.g., MeetingBank) and multi-lingual datasets (e.g., ELITR, FREDSum) is of interest to study our findings across languages. Considering the current evolution of techniques, we further plan to extend the work as a (dialogue) summarization benchmarking pipeline to better capture how upcoming techniques handle the challenges and how well new metrics capture the performance of techniques. We encourage the research community to join our initiative by contributing model outputs and introducing new metrics.

### v2

@@ 4.3 Metrics
We align human annotations with automatic metrics considering count-based, model-based, and QA-based methodologies, chosen based on their prevalent use in meeting and dialogue evaluation Gao and Wan (2022).
Count-based.
@@
Model-based.
BERTScore Zhang et al. (2020b) measures the contextual similarity between generated and reference texts using a pre-trained BERT model, reflecting semantic and syntactic similarity  [^3] We report the rescaled BERTScore-F score:
https://github.com/Tiiiger/bert score/blob/master/journal/rescale baseline.md
. Perplexity (PPL) measures a language model’s uncertainty in predicting words, evaluating the quality and fluency of utterances. We use GPT-2 Radford et al. (2019) as the language model. BLANC Lita et al. (2005) measures how well a generated summary aids a language model in understanding the original document, reflecting informativeness. LENS Maddela et al. (2023) is a trainable metric that assesses the alignment of generated text with human references in content and style. While this metric is also explored for meeting summarization, it is noteworthy that it is based on RoBERTa, adopting its maximum context length of 512 tokens, making it technically less suitable for assessing meeting summaries considering transcripts.
QA-based.
@@
QuestEval Scialom et al. (2021) combines FEQA Durmus et al. (2020) and SummaQA Scialom et al. (2019) scores, using a question-answering model to answer questions formed from the reference text (SummaQA) or the generated summary (FEQA), extracting answers from the opposite source. QuestEval evaluates factuality, coherence, informativeness, and relevance.
Human annotation.
@@
We adapt proven methodologies Zhang et al. (2023) for a thorough annotation process involving four graduate students from diverse backgrounds (i.e., computer science, psychology, communication science), all well-versed in English and familiar with meeting summarization. From the QMSum general-summary test set, we choose 35 general-summary samples, each containing meeting transcripts, gold summaries, and model-generated summaries, resulting in 175 distinct samples for annotation. Annotators identify challenges (Section 3.1) in the transcript and errors Section 3.2 in the generated summaries using yes/no questions such as: "Does the given summary omit crucial information or provide insufficient detail about salient points?" corresponding to the missing information error. The annotators further rate the frequency of challenges and the impact of errors for the LED model using Likert scales from rare occurrence or minimal impact (1) to frequent presence or high impact (5).
To ensure reliability and consistency, we assess inter-annotator agreement using Krippendorff’s alpha, achieving an average of 0.81 (see Tables 2 and 3). A preliminary pilot test serves as annotator training and refinement of guidelines. Regular review meetings are held to maintain consistency. Annotators also highlight summary segments with errors. Discrepancies are discussed, and an expert annotator is available to discuss complex issues.
A full presentation of the annotation process is stated in Appendix A. Details on annotated labels are shown in Appendix B.
@@ 5.1 Linking challenges and error types
Errors Encoder-Decoder Decoder-only
@@
incorrect reasoning speaker dynamics, implicit context
@@
coreference*, contextual turns
@@
Table 4: The linkage between challenges and errors. No asterisk indicates low correlation, * signifies mid correlation (p $\leq$ 0.05), and ** denotes high correlation (p $\leq$ 0.01).
We analyze the relationship between challenges and frequently observed errors for encoder-decoder and decoder-only architectures using Point-biserial correlation based on human annotations from Section 4.3. Table 4 presents the results.
Encoder-decoder models exhibit strong links between incoherence, structural disorganization, and redundancy errors and challenges like spoken language and speaker dynamics, suggesting struggles with maintaining coherence. Conversely, wrong references, linguistic inaccuracy, and hallucination errors show limited associations with the examined challenges. Specifically, the limited association of hallucination reinforces current understandings, suggesting that the underlying causes remain elusive Maynez et al. (2020). Notably, the implicit context challenge works as a proxy for hallucination. The minimal correlation of the wrong reference error, which rarely occurs due to the sentence structure in generated summaries, and the linguistic inaccuracy error indicate that challenges such as coreference are well-handled by these models.
Regarding specific challenges, the discourse structure challenge correlates slightly with the missing information error, indicating occasionally missed details within distinct phases. The low information density challenge weakly correlates with the incorrect reasoning error, revealing difficulties in extracting salient details. The contextual turn-taking challenge aligns with redundancy and incoherence errors, suggesting issues in capturing dynamics on a granular level within the different meeting phases.
Decoder-only models show different patterns, with incorrect reasoning being the rarest and weakest correlated error. Wrong references, redundancy, and linguistic inaccuracy errors are most prevalent, aligning with low information and spoken language challenges. For LLMs, redundancies manifest as repetitive introductory sentences, while linguistic inaccuracies emerge as contextually ambiguous terms. Missing information and structural disorganization errors are equally prominent and strongly correlated, linking to the coreference challenge and emphasizing models’ tendencies to list topics without proper context. Some negative correlations suggest specific challenges might decrease the likelihood of certain errors, such as for low information density challenge and missing information error, warranting further exploration.
@@ 5.2 Correlation of automatic metrics and human annotation
Though numerous and frequently applied, existing automatic metrics are not tailored to the intricacies of meeting summarization. We analyze their reactions to the diverse error types observable in meeting summaries by answering the following three research questions.
@@
Table 5: Point-biserial correlation between automatic metrics and annotated errors on samples generated by LED. * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01. The top row shows abbreviated error types from Section 3.2 in order. A negative correlation indicates worsening metric scores with increasing error occurrence. The three highest absolute values are highlighted in bold. We present F-scores (F) for BERTScore and QuestEval, with BERTScores being further rescaled.
@@
Table 5 shows Point-biserial correlations between automatic metrics and expert-annotated errors for LED model-generated summaries. Our analysis reveals that no metric consistently correlates highly with all error types, underscoring the complexity of meeting summarization and the absence of a universal metric that captures human judgment well Gao and Wan (2022). However, some metrics show trends aligning with human judgment through significant negative scores. Though not designed for structural coherence, several metrics show significant negative correlations with structural disorganization errors, indicating that temporal and logical disorganization breaks n-gram sequences and semantic flow. ROUGE-1 exhibits a more significant score than ROUGE-2 and ROUGE-L, aligning with observations in dialogue summarization Gao and Wan (2022). In particular, the gaps, disjoint narratives, and incorrect statements from errors such as missing information, incoherence, and incorrect reasoning influence metric scores by not aligning with reference n-grams or shifting meaning Lin (2004). Redundancy, wrong references, and hallucination errors remain less frequently detected, as summaries with these errors can still closely match the reference in terms of n-gram overlap and semantic similarity.
@@
As expected, count-based metrics are responsive to missing information Lin (2004), with ROUGE-1 correlating significantly with information omissions Gao and Wan (2022) and BLEU identifying summaries with key content omissions. Count-based metrics significantly correlate with structural disorganization errors though not designed to detect this error, possibly because of disrupted n-gram sequences due to reorganized content Banerjee and Lavie (2005). ROUGE and Meteor are potential indicators of incorrect reasoning and incoherence.
@@
Among model-based metrics, BERTScore exhibits sensitivity to missing information, reflecting semantic and contextual differences between candidate and gold summaries Zhang et al. (2020b). However, its correlation with structural disorganization is less direct, hinting at the influence of sentence alignment and structural coherence. BLANC, designed for evaluating coherence and fluency, correlates as expected negatively with incoherence. LENS effectively captures redundancy errors. Model-based metrics present milder correlations, possibly due to discrepancies between training data and meeting contexts Gao and Wan (2022). While these metrics seem to not align as well with human judgment as count-based metrics, they offer insights into errors not captured by count-based metrics.
@@
The correlation table shows that no single metric predominantly captures all error types, with most correlations being weak to moderate. Examining Table 5, a combination of metrics like ROUGE, BLANC, and LENS could serve as a proxy for various errors, but this approach warrants additional evaluation and score weighting.
@@
Table 6: Spearman correlation between metrics and annotated error impacts, using summaries generated by LED as base. * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01. The top row lists abbreviated error types from Section 3.2 in sequence. A negative correlation implies declining metric scores with rising error instances. The three most pronounced absolute values are emphasized in bold. F-scores (F) are given for BERTScore and QuestEval, with BERTScores being further rescaled.
@@
We categorize "masking" as indifference to an error (near-zero correlation) or positive reinforcement of a mistake (positive correlation). Count-based metrics predominantly show near-zero or negative correlations with errors, indicating they might not sufficiently penalize specific mistakes. This behavior is expected Saadany and Orasan (2021); Akter et al. (2022), especially for errors these metrics were not explicitly designed to detect or for which they can only act as a proxy.
From Table 5 using summaries generated by LED as a base, we observe that some metrics show their known struggle with specific error types through near-zero scores (e.g., BERTScore and linguistic inaccuracy Hanna and Bojar (2021)), model-based metrics occasionally manifest positive correlations. Perplexity favors incorrect references if they preserve linguistic coherence and fluency, aligning with the language model’s expectations and yielding a lower (better) score. LENS appears to struggle with capturing broader logical and temporal structures. Despite being designed for text simplification, it significantly correlates with the structural disorganization error and primarily emphasizes word-level and intra-sentence simplification Maddela et al. (2023). While QuestEval and BETScore do not exhibit significant correlations, they show noteworthy trends: QuestEval may reward errors like missing information and hallucination, a counterintuitive outcome given its focus on factual accuracy. The finding could indicate that hallucinated details likely bypass metric detection due to generated questions lacking the rigor to spot them, which can be linked to the gap between pre-training and meeting data Gao and Wan (2022). BERTScore, emphasizing semantic and syntactic similarities, may reward redundant yet correct details Hanna and Bojar (2021).
@@
To identify the relationship between metric scores and error severity, we analyze how these scores fluctuate by the impact of errors. Table 6 shows the Spearman correlation trends, using summaries generated by LED as a base. Across different models and error severities, most correlations are negligible to weak, and many lack statistical significance. This observation emphasizes the limitations of current metrics in discerning error severity in meeting summarization.
BLEU and QuestEval display significant positive correlations with hallucination errors, underscoring their vulnerability to hallucinated content. While metrics such as BERTScore and LENS show sensitivity to missing information in RQ1, their precision in assessing error severity is limited. This limitation may stem from the gap between training data and meeting transcripts for the metrics utilizing language models, as the meeting transcript was provided as part of the input Gao and Wan (2022). LENS, closely aligned with text simplification, penalizes redundancy. These observations underscore that some metrics are intrinsically responsive to specific errors, even if not explicitly designed for them. Perplexity does not exhibit significant correlations.
@@ 6 Related Work
Research on meeting summarization typically addresses challenges and error types in isolation, leaving their interrelation unexplored. For dialogue summarization, challenges are collected Chen and Yang (2020) but lack the detail required for meeting contexts. Previous works by Tang et al. (2022); Wang et al. (2022) lay the groundwork for our error typologies, enhanced with insights from diverse studies and annotator feedback. Chen and Yang (2020) correlate challenges and errors in dialogue summarization, but their findings do not entirely transfer to English abstractive meeting summarization, though their annotation approach informs ours. Automatic metrics for meeting summaries are underexplored, but similar studies exist for text summarization using the CNN/Dailymail dataset Fabbri et al. (2021) and dialogue summarization via the SAMSum dataset Gao and Wan (2022). We build upon these works by selectively adopting their metrics and augmenting them with measures commonly reported in recent meetings and dialogue summarization studies.
@@ 7 Conclusion
We developed a resource suite for meeting summarization evaluation, covering domain-specific challenges, linked error types, predictive correlations for model architectures with known challenges, and expert annotations of the QMSum subset. Our analysis highlighted the limitations and misalignments of current metrics with human judgment in discerning error nuances. Metrics that work well for other summarization tasks either did not react to errors or cannot reflect the impact on quality in their scores. A composite metric may be more effective, but its formulation requires further research. Recent advancements highlighted the potential of LLM-based metrics. Utilizing an LLM prompted with our detailed annotator guidelines and supplemented by examples presents a promising approach for detecting complex errors like structural disorganization and incorrect reasoning. Given the current evolution of techniques, we plan to extend the work as a (dialogue) summarization benchmarking pipeline to capture better how upcoming techniques handle challenges and how well new metrics capture their performance. We encourage the research community to contribute model outputs and introduce new metrics to this initiative.
Acknowledgements
This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Frederic Kirstein was supported by the Mercedes-Benz AG Research and Development.